# What You Trust Is Insecure: Demystifying How Developers (Mis)Use Trusted Execution Environments in Practice

### v2

@@ I Introduction
To safeguard critical computations from potentially compromised software or operating systems [zheng2021survey], major hardware vendors such as ARM [mcgillion2015open] and Intel [costan2016intel] have introduced Trusted Execution Environments (TEEs) into their processor architectures [li2023survey]. A TEE is an isolated region of the CPU and memory that protects sensitive code and data from untrusted software components through hardware-enforced isolation mechanisms such as encryption and access control [arfaoui2014trusted, zheng2021survey, jauernig2020trusted]. As illustrated in Figure 1, these mechanisms typically partition the system into two domains: the normal world, which hosts untrusted software and system components, and the TEE, which executes trusted code and manages sensitive data. This isolation ensures that critical computations executed within TEEs remain confidential and tamper-resistant, thereby meeting the increasing demand for trustworthy computing across various domains such as mobile payment [wang2020building, zheng2016trustpay], cryptocurrency wallets [gentilal2017trustzone], and remote attestation [10628729]. According to GlobalPlatform’s 2024 Annual Report [globalplatformAnnualReport], billions of devices worldwide, from smartphones to cloud servers, now incorporate TEEs, and the global TEE market is projected to reach \$22.3 billion by 2033 [growthmarketreportsTrustedExecution].
@@
Hardware vendors, who are the primary implementers of TEEs, typically disclose only limited details of their internal designs to software developers. Instead, they provide software development kits (SDKs), such as the Intel Software Guard Extensions (SGX) SDK [intel_sgx_sdk] and the Open Portable (OP)-TEE framework [optee_docs], to standardize and facilitate TEE integration in user applications. However, developing TEE-based software presents unique challenges, particularly in managing encrypted memory and controlling communication between trusted and untrusted environments [li2019teev, paju2023sok], which demands considerable security expertise. While prior studies have shown that developers often exhibit insecure practices in cryptographic and other security-critical contexts, such as misusing Application Programming Interfaces (APIs) [nadi2016jumping, arzt2015towards, mousavi2025detecting, galappaththi2024empirical], little is known about how such issues manifest in the uniquely complex setting of TEE development. Existing studies have primarily focused on analyzing TEE-related mobile binaries [bove2024large] or systematizing academic literature on TEE use cases [paju2023sok], leaving a gap for an empirical investigation that addresses a key question: Do developers adhere to the intended usage patterns of official TEE SDKs, or do practical constraints cause them to deviate from recommended practices?
To bridge this gap, we present the first large-scale quantitative and qualitative investigation of how TEEs are used in real-world software projects. Our study focuses on the two most widely-adopted TEEs, Intel SGX [costan2016intel] and ARM TrustZone [mcgillion2015open], which together account for the majority of publicly available TEE development efforts [li2023survey]. Using the GitHub Search API with TEE-related keywords, we followed prior work [humbatova2020taxonomy, borges2016understanding] to iteratively refine our search queries by adjusting repository size filters, thereby maximizing the coverage of relevant projects. This process yielded 1,547 open-source repositories created or updated between January 2011 and May 2025. After excluding tutorial and documentation-only repositories, as well as unpopular ones with fewer than 10 stars, we manually inspected each repository to confirm the presence of actual TEE implementations. The resulting curated dataset consists of 241 repositories selected for detailed analysis (see Section III). We then combine manual annotation and static analysis to answer the following research questions (RQs):
@@
To understand where TEEs are applied in practice, we categorized all 241 repositories into 8 primary application domains (e.g., AI, blockchain, or secure storage) by analyzing TEE-specific API calls using our customized static analysis scripts. The results show that TEEs are increasingly adopted across diverse domains, yet this trend diverges from prior academic hotspots that emphasize security-critical blockchain and cryptographic systems [gentilal2017trustzone, li2023survey, liu2022extending, zheng2016trustpay, zheng2021survey]. In practice, Internet-of-Things (IoT) device security dominates (30%), whereas blockchain-related projects account for only 7%. Interestingly, AI model protection [mo2021ppfl, mo2020darknetz] has emerged as a fast-growing category (12%) with many projects appearing in recent years, signaling a shift toward protecting machine learning models and data within TEEs and aligning with the broader trend of secure AI development [zhang2022teeslice, lo2023trustworthy].
We also categorized the projects based on the functionality implemented or migrated into TEEs. Cryptographic operations (68.2%) represent the most prevalent functionality implemented within TEEs, which is expected. However, many projects (35.6% of Intel SGX and 27.1% of ARM TrustZone projects) reimplement standard cryptographic operations like encryption algorithms, instead of relying on official SDK APIs. Such reimplementations can introduce logical flaws and side-channel vulnerabilities [ma2025diting, yuan2024hypertheft, joy2024physicalsoftwarebasedfault], while also revealing the limited usability of existing SDKs, where unclear documentation and opaque technical details often lead developers to bypass safe and off-the-shelf APIs. Furthermore, our manual inspection of the source code revealed that 25.3% (61 of 241) of the analyzed projects exhibit insecure coding practices, such as missing input validation or hardcoded secrets. These findings suggest that maintaining sound security principles within the uniquely complex TEE environments remains challenging, and even security-conscious developers may turn TEEs into attack surfaces rather than trusted boundaries. We therefore discuss implications for TEE toolchain designers and the broader secure-systems community to better align TEE capabilities and SDK design with real-world developer practices and to foster safer and more accessible TEE development.
@@ II-A Trusted Execution Environments
Trusted Execution Environments protect sensitive code and data from untrusted operating systems and privileged software through hardware-level isolation mechanisms such as memory encryption and access control [zheng2021survey, li2023survey]. The two most widely-deployed TEEs are Intel SGX [costan2016intel] and ARM TrustZone [mcgillion2015open], designed for the Intel x86 and ARM architectures, respectively. Although they differ in implementation details, both share a common design principle of partitioning the system into two domains: the normal world, which runs untrusted applications, and the TEE, which executes trusted code (called an enclave in Intel SGX and a secure world in TrustZone). In TrustZone, these trusted components are also referred to as Trusted Applications (TAs). Both platforms provide SDKs that enable developers to bridge the two domains with minimal effort. Listing 1 illustrates a simplified interaction, where an untrusted application in the normal world invokes a secure hash function implemented inside the TEE via a vendor-provided API. In this example, developers write only a few lines of code to call the secure function, while the underlying SDK automatically manages cross-world communication.
@@
Recent years have witnessed a growing number of TEE applications, spanning secure payment systems [wang2020building, zheng2016trustpay], blockchain protocols [li2019teev, liu2022extending], and privacy-preserving AI inference [mo2020darknetz, mo2021ppfl, yuan2024hypertheft]. However, the security guarantees of TEEs ultimately depend on how developers design, implement, and interact with enclave or TA code. In practice, the SDKs’ abstraction of low-level operations and the limited disclosure of implementation details may cause developers to misuse TEEs or even introduce new vulnerabilities [ma2025diting, cerdeira2020sok, joy2024physicalsoftwarebasedfault], motivating our empirical study of real-world TEE usage patterns and pitfalls.
@@ II-B Research Landscape on TEE Usability and Adoption
The main body of studies on TEEs can be broadly divided into two categories: system abstractions that simplify TEE development and domain-specific applications that leverage TEE capabilities. The first line of work focuses on building frameworks that abstract low-level TEE operations [shinde2017panoply, tsai2017graphene, paju2023sok]. Representative examples include Graphene-SGX [tsai2017graphene], which enables Linux-compatible system calls inside enclaves; SCONE [arnautov2016scone], which supports containerized enclave execution with automatic encryption and attestation; and Keystone [lee2020keystone], which extends TEE support to RISC-V platforms with customizable enclaves. While these systems lower the entry barrier for developers, they also introduce additional complexity and performance overhead, and none have achieved widespread adoption in practice compared to the official SDKs. The second line of work explores domain-specific use cases that leverage TEE capabilities, such as secure data analytics [yang2023vedb], key management [gentilal2017trustzone, han2025autotee], secure networking [li2019teev, liu2022extending], privacy-preserving machine learning [mo2020darknetz, mo2021ppfl, yuan2024hypertheft], and more [kim2019shieldstore, yang2023vedb]. These studies demonstrate the feasibility of secure computation within TEEs across various domains but often rely on handcrafted or highly specialized implementations that may not reflect how developers use TEEs in everyday practice.
Complementing these system-building efforts, several surveys and systematization papers have aimed to capture the broader TEE research landscape. Paju et al. [paju2023sok] provide a comprehensive literature review that categorizes use cases, SDKs, and usability challenges. Although informative, their work focuses on the academic state of TEE research rather than real-world development practices. Bove et al. [bove2024large] analyze TEE usage in Android applications and find that direct API usage is rare; however, their analysis targets mobile binaries rather than open-source projects with actual implementations, still leaving a gap in understanding developer practices.
Prior studies in the software engineering and security communities have shown that even experienced developers frequently misuse security-critical APIs [fahl2012eve, nadi2016jumping, meng2018secure, wijayarathna2019using, gorski2016towards]. However, development within TEEs, which may be particularly error-prone due to unique challenges such as cross-world communication and encrypted memory management, remains largely understudied. This gap motivates our empirical study to provide a more comprehensive understanding of real-world TEE development practices.
@@ III Data Collection & Labeling
Here we describe how we collect, filter, and label TEE-related projects to form the basis of our empirical analysis. The scope of this study is defined along two dimensions: the targeted TEE platforms and the sources of analyzed projects. As discussed in Section II-A, we focus on the two most widely-adopted TEEs, Intel SGX [costan2016intel] and ARM TrustZone [mcgillion2015open], which together dominate real-world deployments. Other TEEs, such as RISC-V Keystone [lee2020keystone], are excluded due to their limited adoption and the lack of publicly available implementations. We specifically target open-source projects hosted on GitHub, the largest collaborative development platform and a rich source of real-world software artifacts [cosentino2016findings]. No restrictions are imposed on programming languages or project timelines, allowing us to capture a comprehensive snapshot of TEE usage across diverse domains and development practices.
@@
The query was executed using the GitHub Search API. Since the GitHub Search API returns at most 1,000 results per query, we adopted an iterative search strategy following the approach of Humbatova et al. [humbatova2020taxonomy]. Specifically, we varied the size:min..max parameter across multiple queries to circumvent this limitation and capture a broader range of repositories. Each query covered a distinct file size range from 0 to 384,000 bytes, which is the maximum allowed by GitHub, with a step size of 250 bytes to ensure comprehensive coverage of relevant projects. Using this strategy, we retrieved 3,452 files related to OP-TEE, mapped to 875 unique repositories, and 2,438 files for Intel SGX, mapped to 672 unique repositories. Each file was associated with its corresponding repository through metadata, and results from all query sources were merged to form a unified dataset. The resulting repositories were created or last updated between January 2011 and May 2025, covering more than 14 years of TEE development and yielding a total of 1,547 unique repositories.
@@
  • Popularity Filter: We retained only repositories with at least 10 GitHub stars. This threshold helps remove inactive or trivial projects while preserving those that have received community attention or sustained maintenance. Prior studies have shown that popularity metrics such as stars are effective indicators of project maturity and relevance [cosentino2016findings, lyu2025my]. This step removed a substantial number of 1,250 repositories.
@@
Following the automated screening, we manually reviewed each repository to confirm the presence of authentic TEE implementations. In contrast to the lightweight check in Step 2, which considered only non-code information such as GitHub stars and project descriptions, this verification involved an in-depth inspection of the source code to determine whether TEE APIs were functionally integrated into the application’s core logic. This process involved some subjective judgment, particularly for repositories with complex code that required domain expertise to assess genuine TEE usage. To mitigate potential bias, each repository was independently reviewed by two authors to ensure consistency in labeling. The two authors achieved an “almost perfect” level of agreement [viera2005understanding], with a Cohen’s $\kappa$ of 0.92. Disagreements arose in only 8 cases and were resolved through discussion and consensus.
@@ IV-A RQ1: In which domains are TEEs adopted in practice?
Although prior studies have identified TEEs as key enablers of secure computing in sectors such as mobile payment and blockchain [gentilal2017trustzone, liu2022extending], their adoption in real-world applications remains insufficiently understood. Industrial usage may have shifted toward emerging areas, diverging from the traditional security-critical domains emphasized in academic research. By systematically mapping the domains of TEE usage, we aim to uncover where TEEs are truly applied in practice and how these trends have evolved, thereby guiding future academic studies toward greater practical relevance.
@@
To derive a comprehensive taxonomy of TEE application domains, we adopted the open card sorting method [begel2014analyze, cruzes2011recommended, stol2016grounded, yang2023users], a qualitative technique widely used for categorizing software artifacts when predefined categories are unavailable. This approach allows categories to emerge through iterative grouping and refinement, which suits our context where TEE projects’ application domains are diverse and domain boundaries are often ambiguous. In our card sorting, each project was assigned to the domain that best represents its core functionality. For example, a repository implementing a secure communication protocol for IoT devices was classified under IoT Device Security but not under Network Security, even if it also involves network encryption components. Using the open card sorting method and our manual inspection of TEE usage domains, we performed categorization in three steps:
@@
  • Blockchain and Cryptocurrency Security (7%): Projects leveraging TEEs for blockchain transaction integrity and privacy, such as confidential smart contracts [liu2025towards].
  • Network Security and Encrypted Communication (17%): Projects securing network data transmission using TEEs, including Transport Layer Security (TLS) endpoints [apostolopoulos1999transport] and encrypted communication channels.
@@
Key Trends and Insights. The domain distribution of TEE projects reveals several notable patterns. First, although prior research has largely emphasized blockchain and cryptographic systems as primary TEE applications [gentilal2017trustzone, zheng2021survey, li2023survey, zheng2016trustpay], these domains together account for only 7% of real-world projects. In contrast, IoT Device Security dominates (30%), where TEEs are primarily used to ensure device integrity and enable secure updates. Second, the prominence of Privacy-Preserving Computing (19%) and Network Security & Encrypted Communication (17%) indicates that developers increasingly employ TEEs not only for “isolation” but also to build secure computation and communication services within TEEs. This shift reflects the growing maturity of TEEs in supporting complex software stacks and the increasing demand for data protection driven by real-world regulatory pressures (e.g., GDPR [zhang2023pacta, istvan2020towards]).
However, an interesting observation is the growing adoption of AI Model Protection (12%), a relatively new domain that signifies a meaningful convergence between academic research on confidential AI [mo2020darknetz, mo2021ppfl, yuan2024hypertheft] and practical developer needs. Prior research works like DarkneTZ [mo2020darknetz] and PPFL [mo2021ppfl] were among the first to demonstrate the feasibility of embedding AI inference within TEEs to prevent model theft and data leakage. The emergence of this domain in our dataset (12%) suggests that privacy risks initially highlighted by academia are now translating into practical developer needs for securing machine learning workloads. We also observe that U-Boot & Secure Boot projects (4%) are less common, given TEEs’ original design goal of ensuring platform integrity. This trend may reflect the increasing complexity of modern systems, where secure boot is often only one component of a broader security architecture, and developers may prioritize higher-level applications over low-level boot integrity.
We also analyzed the temporal evolution of TEE usage across these domains. Figure 2 shows the cumulative growth of projects in each domain from 2011 to 2025. Early projects (2011–2016) primarily focused on IoT Device Security and Secure Boot, reflecting TEE’s initial role in firmware protection. After that, adoption broadened to Privacy-Preserving Computing and Network Security, enabled by the maturation of TEE SDKs and growing interest in encrypted communication and attestation. After 2019, AI Model Protection rapidly emerged, coinciding with the rise of confidential AI [mo2021ppfl], and has recently become one of the most popular application domains. Overall, this trend illustrates a shift from low-level device protection toward data- and AI model–centric confidentiality, marking the evolution of TEEs from hardware isolation mechanisms to a general-purpose foundation for secure computation.

### v3

@@ I Introduction
To safeguard critical computations from potentially compromised software or operating systems [73], major hardware vendors such as ARM [43] and Intel [14] have introduced Trusted Execution Environments (TEEs) into their processor architectures [35]. A TEE is an isolated region of the CPU and memory that protects sensitive code and data from untrusted software components through hardware-enforced isolation mechanisms such as encryption and access control [3, 73, 27]. As illustrated in Figure 1, these mechanisms typically partition the system into two domains: the normal world, which hosts untrusted software and system components, and the TEE, which executes trusted code and manages sensitive data. This isolation ensures that critical computations executed within TEEs remain confidential and tamper-resistant, thereby meeting the increasing demand for trustworthy computing across various domains such as mobile payment [65, 74], cryptocurrency wallets [20], and remote attestation [67]. According to GlobalPlatform’s 2024 Annual Report [21], billions of devices worldwide, from smartphones to cloud servers, now incorporate TEEs, and the global TEE market is projected to reach \$22.3 billion by 2033 [52].
@@
Hardware vendors, who are the primary implementers of TEEs, typically disclose only limited details of their internal designs to software developers. Instead, they provide software development kits (SDKs), such as the Intel Software Guard Extensions (SGX) SDK [25] and the Open Portable (OP)-TEE framework [60], to standardize and facilitate TEE integration in user applications. However, developing TEE-based software presents unique challenges, particularly in managing encrypted memory and controlling communication between trusted and untrusted environments [34, 51], which demands considerable security expertise. While prior studies have shown that developers often exhibit insecure practices in cryptographic and other security-critical contexts, such as misusing Application Programming Interfaces (APIs) [48, 5, 47, 19], little is known about how such issues manifest in the uniquely complex setting of TEE development. Existing studies have primarily focused on analyzing TEE-related mobile binaries [9] or systematizing academic literature on TEE use cases [51], leaving a gap for an empirical investigation that addresses a key question: Do developers adhere to the intended usage patterns of official TEE SDKs, or do practical constraints cause them to deviate from recommended practices?
To bridge this gap, we present the first large-scale quantitative and qualitative investigation of how TEEs are used in real-world software projects. Our study focuses on the two most widely-adopted TEEs, Intel SGX [14] and ARM TrustZone [43], which together account for the majority of publicly available TEE development efforts [35]. Using the GitHub Search API with TEE-related keywords, we followed prior work [24, 8] to iteratively refine our search queries by adjusting repository size filters, thereby maximizing the coverage of relevant projects. This process yielded 1,547 open-source repositories created or updated between January 2011 and May 2025. After excluding tutorial and documentation-only repositories, as well as unpopular ones with fewer than 10 stars, we manually inspected each repository to confirm the presence of actual TEE implementations. The resulting curated dataset consists of 241 repositories selected for detailed analysis (see Section III). We then combine manual annotation and static analysis to answer the following research questions (RQs):
@@
To understand where TEEs are applied in practice, we categorized all 241 repositories into 8 primary application domains (e.g., AI, blockchain, or secure storage) by analyzing TEE-specific API calls using our customized static analysis scripts. The results show that TEEs are increasingly adopted across diverse domains, yet this trend diverges from prior academic hotspots that emphasize security-critical blockchain and cryptographic systems [20, 35, 36, 74, 73]. In practice, Internet-of-Things (IoT) device security dominates (30%), whereas blockchain-related projects account for only 7%. Interestingly, AI model protection [45, 46] has emerged as a fast-growing category (12%) with many projects appearing in recent years, signaling a shift toward protecting machine learning models and data within TEEs and aligning with the broader trend of secure AI development [72, 38].
We also categorized the projects based on the functionality implemented or migrated into TEEs. Cryptographic operations (68.2%) represent the most prevalent functionality implemented within TEEs, which is expected. However, many projects (35.6% of Intel SGX and 27.1% of ARM TrustZone projects) reimplement standard cryptographic operations like encryption algorithms, instead of relying on official SDK APIs. Such reimplementations can introduce logical flaws and side-channel vulnerabilities [40, 70, 28], while also revealing the limited usability of existing SDKs, where unclear documentation and opaque technical details often lead developers to bypass safe and off-the-shelf APIs. Furthermore, our manual inspection of the source code revealed that 25.3% (61 of 241) of the analyzed projects exhibit insecure coding practices, such as missing input validation or hardcoded secrets. These findings suggest that maintaining sound security principles within the uniquely complex TEE environments remains challenging, and even security-conscious developers may turn TEEs into attack surfaces rather than trusted boundaries. We therefore discuss implications for TEE toolchain designers and the broader secure-systems community to better align TEE capabilities and SDK design with real-world developer practices and to foster safer and more accessible TEE development.
@@ II-A Trusted Execution Environments
Trusted Execution Environments protect sensitive code and data from untrusted operating systems and privileged software through hardware-level isolation mechanisms such as memory encryption and access control [73, 35]. The two most widely-deployed TEEs are Intel SGX [14] and ARM TrustZone [43], designed for the Intel x86 and ARM architectures, respectively. Although they differ in implementation details, both share a common design principle of partitioning the system into two domains: the normal world, which runs untrusted applications, and the TEE, which executes trusted code (called an enclave in Intel SGX and a secure world in TrustZone). In TrustZone, these trusted components are also referred to as Trusted Applications (TAs). Both platforms provide SDKs that enable developers to bridge the two domains with minimal effort. Listing 1 illustrates a simplified interaction, where an untrusted application in the normal world invokes a secure hash function implemented inside the TEE via a vendor-provided API. In this example, developers write only a few lines of code to call the secure function, while the underlying SDK automatically manages cross-world communication.
@@
Recent years have witnessed a growing number of TEE applications, spanning secure payment systems [65, 74], blockchain protocols [34, 36], and privacy-preserving AI inference [46, 45, 70]. However, the security guarantees of TEEs ultimately depend on how developers design, implement, and interact with enclave or TA code. In practice, the SDKs’ abstraction of low-level operations and the limited disclosure of implementation details may cause developers to misuse TEEs or even introduce new vulnerabilities [40, 10, 28], motivating our empirical study of real-world TEE usage patterns and pitfalls.
@@ II-B Research Landscape on TEE Usability and Adoption
The main body of studies on TEEs can be broadly divided into two categories: system abstractions that simplify TEE development and domain-specific applications that leverage TEE capabilities. The first line of work focuses on building frameworks that abstract low-level TEE operations [54, 61, 51]. Representative examples include Graphene-SGX [61], which enables Linux-compatible system calls inside enclaves; SCONE [4], which supports containerized enclave execution with automatic encryption and attestation; and Keystone [31], which extends TEE support to RISC-V platforms with customizable enclaves. While these systems lower the entry barrier for developers, they also introduce additional complexity and performance overhead, and none have achieved widespread adoption in practice compared to the official SDKs. The second line of work explores domain-specific use cases that leverage TEE capabilities, such as secure data analytics [68], key management [20, 23], secure networking [34, 36], privacy-preserving machine learning [46, 45, 70], and more [30, 68]. These studies demonstrate the feasibility of secure computation within TEEs across various domains but often rely on handcrafted or highly specialized implementations that may not reflect how developers use TEEs in everyday practice.
Complementing these system-building efforts, several surveys and systematization papers have aimed to capture the broader TEE research landscape. Paju et al. [51] provide a comprehensive literature review that categorizes use cases, SDKs, and usability challenges. Although informative, their work focuses on the academic state of TEE research rather than real-world development practices. Bove et al. [9] analyze TEE usage in Android applications and find that direct API usage is rare; however, their analysis targets mobile binaries rather than open-source projects with actual implementations, still leaving a gap in understanding developer practices.
Prior studies in the software engineering and security communities have shown that even experienced developers frequently misuse security-critical APIs [17, 48, 44, 66, 22]. However, development within TEEs, which may be particularly error-prone due to unique challenges such as cross-world communication and encrypted memory management, remains largely understudied. This gap motivates our empirical study to provide a more comprehensive understanding of real-world TEE development practices.
@@ III Data Collection & Labeling
Here we describe how we collect, filter, and label TEE-related projects to form the basis of our empirical analysis. The scope of this study is defined along two dimensions: the targeted TEE platforms and the sources of analyzed projects. As discussed in Section II-A, we focus on the two most widely-adopted TEEs, Intel SGX [14] and ARM TrustZone [43], which together dominate real-world deployments. Other TEEs, such as RISC-V Keystone [31], are excluded due to their limited adoption and the lack of publicly available implementations. We specifically target open-source projects hosted on GitHub, the largest collaborative development platform and a rich source of real-world software artifacts [13]. No restrictions are imposed on programming languages or project timelines, allowing us to capture a comprehensive snapshot of TEE usage across diverse domains and development practices.
@@
The query was executed using the GitHub Search API. Since the GitHub Search API returns at most 1,000 results per query, we adopted an iterative search strategy following the approach of Humbatova et al. [24]. Specifically, we varied the size:min..max parameter across multiple queries to circumvent this limitation and capture a broader range of repositories. Each query covered a distinct file size range from 0 to 384,000 bytes, which is the maximum allowed by GitHub, with a step size of 250 bytes to ensure comprehensive coverage of relevant projects. Using this strategy, we retrieved 3,452 files related to OP-TEE, mapped to 875 unique repositories, and 2,438 files for Intel SGX, mapped to 672 unique repositories. Each file was associated with its corresponding repository through metadata, and results from all query sources were merged to form a unified dataset. The resulting repositories were created or last updated between January 2011 and May 2025, covering more than 14 years of TEE development and yielding a total of 1,547 unique repositories.
@@
  • Popularity Filter: We retained only repositories with at least 10 GitHub stars. This threshold helps remove inactive or trivial projects while preserving those that have received community attention or sustained maintenance. Prior studies have shown that popularity metrics such as stars are effective indicators of project maturity and relevance [13, 39]. This step removed a substantial number of 1,250 repositories.
@@
Following the automated screening, we manually reviewed each repository to confirm the presence of authentic TEE implementations. In contrast to the lightweight check in Step 2, which considered only non-code information such as GitHub stars and project descriptions, this verification involved an in-depth inspection of the source code to determine whether TEE APIs were functionally integrated into the application’s core logic. This process involved some subjective judgment, particularly for repositories with complex code that required domain expertise to assess genuine TEE usage. To mitigate potential bias, each repository was independently reviewed by two authors to ensure consistency in labeling. The two authors achieved an “almost perfect” level of agreement [63], with a Cohen’s $\kappa$ of 0.92. Disagreements arose in only 8 cases and were resolved through discussion and consensus.
@@ IV-A RQ1: In which domains are TEEs adopted in practice?
Although prior studies have identified TEEs as key enablers of secure computing in sectors such as mobile payment and blockchain [20, 36], their adoption in real-world applications remains insufficiently understood. Industrial usage may have shifted toward emerging areas, diverging from the traditional security-critical domains emphasized in academic research. By systematically mapping the domains of TEE usage, we aim to uncover where TEEs are truly applied in practice and how these trends have evolved, thereby guiding future academic studies toward greater practical relevance.
@@
To derive a comprehensive taxonomy of TEE application domains, we adopted the open card sorting method [6, 15, 57, 69], a qualitative technique widely used for categorizing software artifacts when predefined categories are unavailable. This approach allows categories to emerge through iterative grouping and refinement, which suits our context where TEE projects’ application domains are diverse and domain boundaries are often ambiguous. In our card sorting, each project was assigned to the domain that best represents its core functionality. For example, a repository implementing a secure communication protocol for IoT devices was classified under IoT Device Security but not under Network Security, even if it also involves network encryption components. Using the open card sorting method and our manual inspection of TEE usage domains, we performed categorization in three steps:
@@
  • Blockchain and Cryptocurrency Security (7%): Projects leveraging TEEs for blockchain transaction integrity and privacy, such as confidential smart contracts [37].
  • Network Security and Encrypted Communication (17%): Projects securing network data transmission using TEEs, including Transport Layer Security (TLS) endpoints [2] and encrypted communication channels.
@@
Key Trends and Insights. The domain distribution of TEE projects reveals several notable patterns. First, although prior research has largely emphasized blockchain and cryptographic systems as primary TEE applications [20, 73, 35, 74], these domains together account for only 7% of real-world projects. In contrast, IoT Device Security dominates (30%), where TEEs are primarily used to ensure device integrity and enable secure updates. Second, the prominence of Privacy-Preserving Computing (19%) and Network Security & Encrypted Communication (17%) indicates that developers increasingly employ TEEs not only for “isolation” but also to build secure computation and communication services within TEEs. This shift reflects the growing maturity of TEEs in supporting complex software stacks and the increasing demand for data protection driven by real-world regulatory pressures (e.g., GDPR [71, 26]).
However, an interesting observation is the growing adoption of AI Model Protection (12%), a relatively new domain that signifies a meaningful convergence between academic research on confidential AI [46, 45, 70] and practical developer needs. Prior research works like DarkneTZ [46] and PPFL [45] were among the first to demonstrate the feasibility of embedding AI inference within TEEs to prevent model theft and data leakage. The emergence of this domain in our dataset (12%) suggests that privacy risks initially highlighted by academia are now translating into practical developer needs for securing machine learning workloads. We also observe that U-Boot & Secure Boot projects (4%) are less common, given TEEs’ original design goal of ensuring platform integrity. This trend may reflect the increasing complexity of modern systems, where secure boot is often only one component of a broader security architecture, and developers may prioritize higher-level applications over low-level boot integrity.
We also analyzed the temporal evolution of TEE usage across these domains. Figure 2 shows the cumulative growth of projects in each domain from 2011 to 2025. Early projects (2011–2016) primarily focused on IoT Device Security and Secure Boot, reflecting TEE’s initial role in firmware protection. After that, adoption broadened to Privacy-Preserving Computing and Network Security, enabled by the maturation of TEE SDKs and growing interest in encrypted communication and attestation. After 2019, AI Model Protection rapidly emerged, coinciding with the rise of confidential AI [45], and has recently become one of the most popular application domains. Overall, this trend illustrates a shift from low-level device protection toward data- and AI model–centric confidentiality, marking the evolution of TEEs from hardware isolation mechanisms to a general-purpose foundation for secure computation.